# MetaSym: A Symplectic Meta-learning Framework for Physical Intelligence

### v2

@@ Abstract
Scalable and generalizable physics-aware deep learning has long been considered a significant challenge with various applications across diverse domains ranging from robotics to molecular dynamics. Central to almost all physical systems are symplectic forms—the geometric backbone that underpins fundamental invariants like energy and momentum. In this work, we introduce a novel deep learning framework, MetaSym. In particular, MetaSym combines a strong symplectic inductive bias obtained from a symplectic encoder, and an autoregressive decoder with meta-attention. This principled design ensures that core physical invariants remain intact, while allowing flexible, data-efficient adaptation to system heterogeneities. We benchmark MetaSym with highly varied and realistic datasets, such as a high-dimensional spring-mesh system [1], an open quantum system with dissipation and measurement backaction, and robotics-inspired quadrotor dynamics. Our results demonstrate superior performance in modeling dynamics under few-shot adaptation, outperforming state-of-the-art baselines that use larger models.
@@ 1 Introduction
Learning to predict the dynamics of physical systems is a fundamental challenge in scientific machine learning, with applications ranging from robotics, control, climate science, and quantum computing [2, 3, 4]. Traditional approaches often rely on carefully derived differential equations that embed known conservation laws and geometric properties [5] (e.g., Hamiltonian or Lagrangian mechanics). Although these classical models have been tremendously successful in capturing fundamental physics, they can become unwieldy or intractable when faced with complex real-world phenomena, such as high-dimensional systems, intricate interactions, or partially unknown forces, that defy simple closed-form representations. Recent progress in deep learning has opened new avenues for data-driven modeling of dynamical systems, bypassing the need for complete analytical descriptions [6]. Neural ODE frameworks [7], for instance, reinterpret dynamic evolution as a continuous function learned by a neural network, while operator-learning approaches such as Fourier Neural Operators (FNOs) [8] allow for flexible mappings from initial conditions to solutions of partial differential equations. Despite these advances, deep learning approaches often face two critical challenges:
@@
  • Preserving the underlying physical structure. Standard networks, left unconstrained, may inadvertently violate symplectic forms, conservation of energy, or other geometric constraints intrinsic to physical dynamics [9]. These violations can accumulate over time, producing qualitatively incorrect long-horizon predictions, e.g. spurious energy drift in Hamiltonian systems.
@@
  • Generalizing across related systems. Many real-world applications involve entire families of similar yet distinct systems (e.g., variations of a robotic manipulator differing in load mass or joint friction, or molecular systems differing in exact bonding parameters). Training an entirely separate model for each variant is both data-inefficient and computationally expensive. Without mechanisms to share knowledge, a network trained on one system will typically fail to adapt efficiently to another, even if they exhibit similar physical behaviors. Bridging this sim-2-real gap is crucial for a variety of tasks such as control and real-time prediction [10].
@@
The trade-off between flexibility (i.e., capacity to learn diverse dynamics) and the enforcement of physical constraints can be addressed through specialized architectures that embed geometric priors [9, 11] related to the physical system. Notably in the context of Hamiltonian mechanics, Symplectic Networks (SympNets) incorporate symplectic forms directly into their design, guaranteeing that learned transformations represent canonical transformations [12]. This preserves key invariants such as energy and momentum, mitigating error accumulation in long-horizon forecasts.
@@
However, real-world systems also exhibit heterogeneity: varying parameters, boundary conditions, or even control signals that deviate from conservative dynamics. Thus, meta-learning becomes a natural extension [13]. By training on a set of related systems, meta-learning-based methods (e.g., Model-Agnostic Meta-Learning (MAML), interpretable Meta neural Ordinary Differential Equation (iMODE) and Fast Context Adaptation Via Meta-Learning (CAVIA) [14, 15, 16]) acquire high-level inductive biases that can be quickly adapted to novel systems using limited additional data. Consequently, when one faces a new variant of a familiar system, the network can fine-tune a handful of parameters, rather than retrain from scratch. This provides robust and scalable performance.
@@ 1.1 Contributions
  • Symplectic Encoder for Structure Preservation: We propose a specialized neural encoder (SymplecticEncoder) built on SympNet modules. The inherent structural invariants of the SympNets provide a strong inductive bias to the SymplecticEncoder, while our bi-directional training pipeline enforces Hamiltonian consistency, obtained by the canonical transformations that pertain to different systems. Hence, the SymplecticEncoder’s output conserves key geometric invariants (e.g., energy and momentum), effectively minimizing error accumulation over long-term rollouts with minimal architecture size.
  • Autoregressive Decoder with Meta-Attention for Adaptation: To handle nonconservative forces and variations in system parameters, we introduce ActiveDecoder, a transformer-inspired decoder module equipped with a meta-attention framework. This decoder incorporates control-inputs, external forces, and per-system parameters enabling flexible modeling of real-world effects beyond ideal Hamiltonian dynamics while enabling autoregressive multi-step prediction during inference time.
  • Meta-Learning for Multi-System Generalization: We adopt different meta-learning paradigms for the SymplecticEncoder and ActiveDecoder of our architecture motivated by specific design choices explained in Section 3. This yields a single framework that quickly adapts to new or modified systems with minimal additional data.
By integrating physical constraints and generalizable architectural changes into a novel training pipeline that utilizes meta-learning updates for both our SymplecticEncoder and ActiveDecoder, we benchmark this bespoke smaller architecture against other state-of-the-art deep learning methods, including Dissipative Hamiltonian Neural Networks (DHNNs) [17] and Transformers [18, 19], for modeling various physical systems in both classical and quantum regimes. These systems include a high-dimensional spring mesh system, an open quantum system whose dynamics are highly complex and counterintuitive in the classical regime, and a quadrotor with floating base dynamics. This provides evidence of far reaching implications for a diverse set of physics modeling tasks including challenging tasks like real-time quantum dynamics prediction and simulating a variety of complex dynamics that typically require complex computational methods to solve. To the best of our knowledge, MetaSym is the first bespoke physics-based deep learning model, to adapt and generalize well to both classical and non-unitary quantum dynamics.
@@ 2 Related Work and Background
Physicists have long utilized the Lagrangian and Hamiltonian formalisms of mechanics to study the dynamics of physical systems [20]. Consider $\mathbf{q}=\left(q_{1},\ldots,q_{d}\right)$ that represents the generalized coordinates of an $d$-dimensional configuration space (position), while ${\mathbf{p}}$ represents the corresponding generalized momenta. We can describe the Hamiltonian $H(\mathbf{q},\mathbf{p},t)$ as the total energy of the system. This leads to Hamilton’s equations as follows:
@@
This formalism naturally imbues the phase space $(\mathbf{q},\mathbf{p})\in\mathbb{R}^{2d}$ with geometric structures that are vital to the study of these physical systems. One such structure is the symplectic form ($\omega$) given by, $\omega=\sum_{i=1}^{n}\mathrm{~{}d}p_{i}\wedge\mathrm{~{}d}q_{i}$. Concretely, a map $\Phi_{t}:(\mathbf{q}(0),\mathbf{p}(0))\mapsto(\mathbf{q}(t),\mathbf{p}(t))$ is said to be symplectic if $\Phi_{t}^{*}\omega=\omega$. This implies that the flow in the phase space preserves the volume [21].
@@
Inspired by such geometric formulations of mechanics, recent work such as Hamiltonian Neural Networks (HNNs) and Lagrangian Neural Networks (LNNs) have sought to embed physical priors into deep learning architectures [11, 22]. In particular, SympNets have emerged as structure-preserving neural architectures designed specifically for learning Hamiltonian systems, ensuring that their learned maps are intrinsically symplectic. These architectures admit universal approximation theorems and crucially, this construction does not require solving ODEs or differentiating a candidate Hamiltonian during training and inference times, which often leads to more efficient optimization compared to other architectures such as HNNs or LNNs. The collection of all SympNets forms a group under composition, ensuring that every learned map is invertible with a closed-form inverse. However, due to the fundamental nature of Hamiltonian systems and symplectic forms, SympNets and similar architectures fail to generalize to dissipative systems where the symplectic structure is no longer preserved [9, 22, 12]. While there have been attempts to reconcile dissipative dynamics and control inputs to model realistic physical systems by preserving the symplectic framework, such as the Dissipative SymODEN [23, 24], they often suffer from the lack of generalization to different systems [25].
@@
Generalization between different but related physical systems is also vital for deep learning methods to excel at physics modeling. Meta-learning serves as the natural avenue for exploring such strategies. Building on a series of meta-learning strategies [14, 26, 27, 28], the iMODE framework represents a notable advancement in applying deep learning to families of dynamical systems. In this setup, one set of parameters captures the universal dynamics shared across all systems in a family, while another set encodes the idiosyncratic physical parameters that differentiate one system instance from another. However, certain drawbacks still persist. Apart from the lack of physics priors, these existing meta-learning approaches such as iMODE suffer from a lack of scalability [29].
@@
Figure 1: MetaSym integrates a SymplecticEncoder (light-green), ActiveDecoder (light-purple), and ControlNet (pink). The SymplecticEncoder is pre-trained in isolation on conservative state-space data using shared forward/inverse networks receiving $(\mathbf{q}^{(i)}_{t},\mathbf{p}^{(i)}_{t},\Delta t)$ and $(\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1},\Delta t)$ respectively, and thus explicitly enforcing time-reversibility. Subsequently, with the SymplecticEncoder frozen, the ActiveDecoder and ControlNet are then trained autoregressively via teacher forcing, where system-specific adaptation is achieved by fine-tuning the cross-attention’s query/value parameters (purple dots in cross attention) with few-shot gradient steps. During inference, the ControlNet processes a sequence of non-conservative coordinates and control signals $\{\mathbf{\tilde{q}}^{(i)}_{t:T},\mathbf{\tilde{p}}^{(i)}_{t:T},\Delta t,\mathbf{\tilde{u}}^{(i)}_{t:T}\}$, while the SymplecticEncoder projects them onto the conservative manifold and integrates them on it, producing $\mathbf{z}_{c}$. The ActiveDecoder using its cross-attention, perturbs $\mathbf{z}_{c}$ to predict the dynamics of the system for future timesteps, autoregressively. We indicate the next-step predictions of our network with $\left(\mathbf{\hat{q}}^{(i)}_{t+1},\mathbf{\hat{p}}^{(i)}_{t+1}\right)$.
@@ 3 Methods
In this section, we detail our pipeline for learning, adapting, and predicting the dynamics of physical systems using, MetaSym, our structure-preserving neural architecture and meta-learning framework. We describe both the high-level design of our encoder–decoder model and the specialized training procedures implemented via meta-learning. Figure 1 represents the overview of the architecture. Details on the described training and meta-learning pipelines are given in Appendix A, while each design choice is based on ablation studies provided in Appendix D.
@@ 3.1 SymplecticEncoder: Structure Preservation
To ensure that predictions preserve fundamental geometric invariants, the encoder module is implemented as a SymplecticEncoder. Internally, it uses a symplectic neural network (e.g., LASympNet [12]) consisting of sub-layers that update position $\mathbf{q}^{(i)}_{t}$ and momentum $\mathbf{p}^{(i)}_{t}$ for systems $i=1,...,n$ in a manner designed to approximate Hamiltonian flows and preserve symplecticity. Specifically, each sub-layer either performs an “up” or “low” transformation,
[…]
Here, we use $\alpha$ and $\beta$ that can be either linear or activation modules, ensuring we remain in the class of canonical (i.e., symplectic) maps. By stacking multiple up/low blocks, we achieve a deep network $\Phi_{\theta}:(\mathbf{q^{(i)}_{t}},\mathbf{p^{(i)}_{t}},\Delta t)\mapsto(\mathbf{q^{(i)}_{t+1}},\mathbf{p^{(i)}_{t+1}})$. A crucial mathematical property is that the composition of symplectic transformations remains symplectic. Thus, when these modules are stacked together in sequence, regardless of network depth or configuration, the resulting transformation is guaranteed to be symplectic.
A pivotal feature of Hamiltonian dynamics is time reversibility: if we integrate the system forward from $(\mathbf{q^{(i)}_{t}},\mathbf{p^{(i)}_{t})}$ to $(\mathbf{q^{(i)}_{t+1}},\mathbf{p^{(i)}_{t+1}})$ over time $\Delta t$, then integrating backwards over $-\Delta t$ should return the system exactly to $(\mathbf{q^{(i)}_{t}},\mathbf{p^{(i)}_{t}})$. This property lies at the heart of many physical invariants (energy, momentum, etc.) and is crucial for long-horizon stability in forecasting the behavior of such systems.
@@
During training the encoder is provided with a sequence of phase-space points (T-timesteps) and encodes them onto a conservative latent space $\mathbf{z}_{c}=[\mathbf{q_{enc}},\mathbf{p_{enc}}]\in\mathbb{R}^{2d\times T}$ that is ultimately used by the ActiveDecoder in the cross attention mechanism to specifically fine-tune the query and value parameters.
Additionally, to ensure that our encoder can generalize quickly across multiple related but distinct physical systems, we adopt a MAML-style [14] framework. This meta-learning paradigm works particularly well in the SymplecticEncoder’s case, since SympNets represent canonical transformations that are known to have well-defined Hessian [30], which is crucial for methods like MAML that use second-order gradients. For each system $i$ in a mini-batch, we split its trajectory into $\mathcal{I}_{adapt}$ and $\mathcal{I}_{meta}$ sets. During the fast adaptation loop, we optimize the parameters of the encoder $\theta_{SE}$ only on $\mathcal{I}_{adapt}$, by minimizing the mean-squared error between its forward predictions and the ground-truth labels. This process simulates “specializing” the encoder to system $i$’s local dynamics using a simple loss function that avoids gradient-terms that may destabilize the subsequent meta-update of $\theta_{SE}$. For the outer loop, we perform a forward and an inverse pass of the SymplecticEncoder and subsequently we minimize a combined loss represented as,
$\mathcal{L}_{\mathrm{meta}}=\frac{1}{\mathcal{T}_{meta}N_{batch}}$ $\sum_{\begin{subarray}{c}t\in\mathcal{I}_{\mathrm{meta}}\\ i\in N_{batch}\end{subarray}}\Bigl{\|}\Phi_{\theta_{SE}}\bigl{(}\mathbf{q}_{t}^{(i)},\mathbf{p}_{t}^{(i)};\boldsymbol{\theta_{SE}}^{*(i)}\bigr{)}-\Phi_{\theta_{SE}}^{-1}\bigl{(}\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)};\boldsymbol{\theta_{SE}}^{*(i)}\bigr{)}$ (3)
$-[\mathbf{q}_{t+1}^{(i)}-\mathbf{q}_{t}^{(i)},\mathbf{p}_{t+1}^{(i)}-\mathbf{p}_{t}^{(i)}]\Bigr{\|}^{2},$
where $\mathcal{T}_{meta}$ is the total number of timesteps stored in $\mathcal{I}_{meta}$ set and $N_{batch}$ the total number of meta-learned systems in the batch. While the SymplecticEncoder is trained in isolation, we freeze its parameters during the ActiveDecoder’s training.

### v3

@@ Abstract
Scalable and generalizable physics-aware deep learning has long been considered a significant challenge with various applications across diverse domains ranging from robotics to molecular dynamics. Central to almost all physical systems are symplectic forms, the geometric backbone that underpins fundamental invariants like energy and momentum. In this work, we introduce a novel deep-learning framework, MetaSym. Our approach combines a strong symplectic inductive bias obtained from a symplectic encoder, and an autoregressive decoder with meta-attention. This principled design ensures that core physical invariants remain intact, while allowing flexible, data efficient adaptation to system heterogeneities. We benchmark MetaSym with highly varied and realistic datasets, such as a high-dimensional spring-mesh system (Otness et al., 2021), an quantum system with dissipation and measurement backaction, and robotics-inspired quadrotor dynamics. Crucially, we fine-tune and deploy MetaSym on real-world quadrotor data, demonstrating robustness to sensor noise and real-world uncertainty. Across all tasks, MetaSym achieves superior few-shot adaptation and outperforms larger state-of-the-art (SOTA) models.
@@ 1 Introduction
Learning to predict the dynamics of physical systems is a fundamental challenge in scientific machine learning, with applications ranging from robotics, control, climate science, and quantum computing (Ghadami & Epureanu, 2022; Zhang et al., 2024; Alexeev et al., 2024). Traditional approaches often rely on carefully derived differential equations (Papatheodorou et al., 2024) that embed known conservation laws and geometric properties (e.g., Hamiltonian or Lagrangian mechanics). Although these classical models have been tremendously successful in capturing fundamental physics, they can become unwieldy or intractable when faced with complex real-world phenomena, such as high-dimensional systems, intricate interactions, or partially known forces, that defy simple closed-form representations. Recent progress in deep learning has opened new avenues for data-driven modeling of dynamical systems, bypassing the need for complete analytical descriptions (Carleo et al., 2019). Neural ODE frameworks such as Chen et al. (2019), for instance, reinterpret dynamic evolution as a continuous function learned by a neural network, while operator-learning approaches such as Fourier Neural Operators (FNOs) in Li et al. (2021) allow for flexible mappings from initial conditions to solutions of partial differential equations. Despite these advances, deep learning approaches often face two critical challenges:
@@
  • Preserving the underlying physical structure. Standard networks, left unconstrained, may inadvertently violate symplectic forms, conservation of energy, or other geometric constraints intrinsic to physical dynamics (Chen et al., 2020). These violations can accumulate over time, producing qualitatively incorrect long-horizon predictions, e.g. spurious energy drift in Hamiltonian systems.
@@
  • Generalizing across related systems. Many real-world applications involve entire families of similar yet distinct systems (e.g., variations of a robotic manipulator differing in load mass or joint friction, or molecular systems differing in exact bonding parameters). Training an entirely separate model for each variant is both data-inefficient and computationally expensive. Without mechanisms to share knowledge, a network trained on one system will typically fail to adapt efficiently to another, even if they exhibit similar physical behaviors. Bridging this sim-2-real gap is crucial for a variety of tasks such as control and real-time prediction Bai et al. (2024).
@@
The trade-off between flexibility (i.e., capacity to learn diverse dynamics) and the enforcement of physical constraints can be addressed through specialized architectures that embed geometric priors (Chen et al., 2020; Greydanus et al., 2019) related to the physical system. Notably in the context of Hamiltonian mechanics, the Symplectic Networks (SympNets), introduced in Jin et al. (2020), incorporate symplectic forms directly into their design, guaranteeing that learned transformations represent canonical transformations. This preserves key invariants such as energy and momentum, mitigating error accumulation in long-horizon forecasts.
@@
However, real-world systems also exhibit heterogeneity, i.e., varying parameters, boundary conditions, or even control signals that deviate from conservative dynamics. Thus, meta-learning becomes a natural extension (Schorling et al., 2025). By training on a set of related systems, meta-learning-based methods (e.g., Model-Agnostic Meta-Learning (MAML), interpretable Meta neural Ordinary Differential Equation (iMODE) and Fast Context Adaptation Via Meta-Learning (CAVIA) (Finn et al., 2017; Li et al., 2023; Zintgraf et al., 2019b)) acquire high-level inductive biases that can be quickly adapted to novel systems using limited additional data. Consequently, when one faces a new variant of a familiar system, the network can fine-tune a handful of parameters, rather than retrain from scratch. This provides robust and scalable performance.
@@ 1.1 Contributions
  • Symplectic encoder for structure preservation. We propose a specialized neural encoder (SymplecticEncoder) built on SympNet modules. The inherent structural invariants of the SympNets provide a strong inductive bias to the SymplecticEncoder, while our bi-directional training pipeline enforces Hamiltonian consistency, obtained by the canonical transformations that pertain to different systems. Hence, the SymplecticEncoder’s output conserves key geometric invariants (e.g., energy and momentum), effectively minimizing error accumulation over long-term roll-outs with minimal architecture size.
  • Autoregressive decoder with meta-attention for adaptation. To handle non-conservative forces and variations in system parameters, we introduce ActiveDecoder, a transformer-inspired decoder module equipped with a meta-attention framework. This decoder incorporates control inputs, external forces, and per-system parameters enabling flexible modeling of real-world effects beyond ideal Hamiltonian dynamics, while enabling autoregressive multi-step prediction during inference time.
  • Meta-learning for multi-system generalization. We adopt different meta-learning paradigms for the SymplecticEncoder and ActiveDecoder of our architecture motivated by specific design choices explained in Section 3. This yields a single framework that quickly adapts to new or modified systems, even in real-world scenarios.
By integrating physical constraints and generalizable architectural changes into a novel training pipeline that utilizes meta-learning updates for both our SymplecticEncoder and ActiveDecoder, we benchmark this bespoke smaller architecture against other SOTA deep learning methods, including Dissipative Hamiltonian Neural Networks (DHNNs) (Sosanya & Greydanus, 2022) and Transformers (Vaswani, 2017; Geneva & Zabaras, 2022), for modeling various physical systems in both classical and quantum regimes. These systems include a high-dimensional spring-mesh system, an open-quantum system whose dynamics are highly complex and counterintuitive in the classical regime, and a quadrotor with floating-base dynamics. This provides evidence of far reaching implications for a diverse set of physics modeling tasks including challenging tasks like real-time quantum-dynamics prediction and simulating a variety of complex dynamics that typically require complex computational methods to solve. To the best of our knowledge, MetaSym is the first bespoke physics-based deep-learning model to adapt and generalize well to both classical and non-unitary quantum dynamics.
@@ 2 Related Work and Background
Physicists have long utilized the Lagrangian and Hamiltonian formalisms of mechanics to study the dynamics of physical systems (Kibble & Berkshire, 2004). Consider $\mathbf{q}=\left(q_{1},\ldots,q_{d}\right)$ that represents the generalized coordinates of an $d$-dimensional configuration space (position), while $\mathbf{p}$ represents the corresponding generalized momenta. We can describe the Hamiltonian $H(\mathbf{q},\mathbf{p},t)$ as the total energy of the system. This leads to Hamilton’s equations as follows:
@@
This formalism naturally imbues the phase space $(\mathbf{q},\mathbf{p})\in\mathbb{R}^{2d}$ with geometric structures that are vital to the study of these physical systems. One such structure is the symplectic form ($\omega$) given by, $\omega=\sum_{i=1}^{n}\mathrm{~d}p_{i}\wedge\mathrm{~d}q_{i}$. Concretely, a map $\Phi_{t}:(\mathbf{q}(0),\mathbf{p}(0))\mapsto(\mathbf{q}(t),\mathbf{p}(t))$ is said to be symplectic if $\Phi_{t}^{*}\omega=\omega$. This implies that the flow in the phase space preserves the volume (Royer, 1991).
@@
Inspired by such geometric formulations of mechanics, recent work such as Hamiltonian Neural Networks (HNNs) and Lagrangian Neural Networks (LNNs) have sought to embed physical priors into deep-learning architectures (Greydanus et al., 2019; Cranmer et al., 2020). In particular, SympNets have emerged as structure-preserving neural architectures designed specifically for learning Hamiltonian systems, ensuring that their learned maps are intrinsically symplectic. These architectures admit universal approximation theorems and crucially, this construction does not require solving ODEs or differentiating a candidate Hamiltonian during training and inference times, which often leads to more efficient optimization compared to other architectures such as HNNs or LNNs. The collection of all SympNets forms a group under composition, ensuring that every learned map is invertible with a closed-form inverse. However, due to the fundamental nature of Hamiltonian systems and symplectic forms, SympNets and similar architectures fail to generalize to dissipative systems where the symplectic structure is no longer preserved (Chen et al., 2020; Cranmer et al., 2020; Jin et al., 2020). While there have been attempts to reconcile dissipative dynamics and control inputs to model realistic physical systems by preserving the symplectic framework, such as the Dissipative SymODEN (Zhong et al., 2020; 2024), they often suffer from the lack of generalization to different systems (Okamoto & Kojima, 2024).
@@
Generalization between different but related physical systems is also vital for deep-learning methods to excel at physics modeling. Meta-learning serves as the natural avenue for exploring such strategies. Building on a series of meta-learning strategies (Finn et al., 2017; Rajeswaran et al., 2019; Zintgraf et al., 2019a; Nichol, 2018), the iMODE framework represents a notable advancement in applying deep learning to families of dynamical systems. In this setup, one set of parameters captures the universal dynamics shared across all systems in a family, while another set encodes the idiosyncratic physical parameters that differentiate one system instance from another. However, certain drawbacks still persist. Apart from the lack of physics priors, the existing meta-learning approaches, such as iMODE, suffer from a lack of scalability (Choe et al., 2023).
@@
Figure 1: MetaSym integrates a SymplecticEncoder (light-green), ActiveDecoder (light-purple), and ControlNet (pink). The SymplecticEncoder is pre-trained in isolation on conservative state-space data using shared forward/inverse networks receiving $(\mathbf{q}^{(i)}_{t},\mathbf{p}^{(i)}_{t},\Delta t)$ and $(\mathbf{q}^{(i)}_{t+1},\mathbf{p}^{(i)}_{t+1},\Delta t)$ respectively, and thus explicitly enforcing time-reversibility. Subsequently, with the SymplecticEncoder frozen, the ActiveDecoder and ControlNet are then trained autoregressively via teacher forcing, where system-specific adaptation is achieved by fine-tuning the cross-attention’s query/value parameters (purple dots in cross attention) with few-shot gradient steps. During inference, the ControlNet processes a sequence of non-conservative coordinates and control signals $\{\mathbf{\tilde{q}}^{(i)}_{t:T},\mathbf{\tilde{p}}^{(i)}_{t:T},\Delta t,\mathbf{\tilde{u}}^{(i)}_{t:T}\}$, while the SymplecticEncoder projects them onto the conservative manifold and integrates them on it, producing $\mathbf{z}_{c}$. The ActiveDecoder using its cross-attention, perturbs $\mathbf{z}_{c}$ to predict the dynamics of the system for future time-steps, autoregressively. We indicate the next-step predictions of our network with $\left(\mathbf{\hat{q}}^{(i)}_{t+1},\mathbf{\hat{p}}^{(i)}_{t+1}\right)$.
@@ 3 Methods
In this section, we detail our pipeline for learning, adapting, and predicting the dynamics of physical systems using MetaSym, our structure-preserving neural architecture and meta-learning framework. We describe both the high-level design of our encoder–decoder model and the specialized training procedures implemented via meta-learning. Fig. 1 represents the overview of the architecture. Details on the described training and meta-learning pipelines are given in Appendix A, while each design choice is based on ablation studies provided in Appendix D.
@@ 3.1 SymplecticEncoder: Structure Preservation
To ensure that predictions preserve fundamental geometric invariants, the encoder module is implemented as a SymplecticEncoder. Internally, it uses a symplectic neural network (e.g., LASympNet by Jin et al. (2020)) consisting of sub-layers that update position $\mathbf{q}^{(i)}_{t}$ and momentum $\mathbf{p}^{(i)}_{t}$ for systems i=1,,ni=1,...,n in a manner designed to approximate Hamiltonian flows and preserve symplecticity. Specifically, each sub-layer either performs an “up” or “low” transformation,
$\text{(Up)}\;\mathbf{q}^{(i)}_{t}\leftarrow\mathbf{q}^{(i)}_{t}\alpha(\mathbf{p}^{(i)}_{t})\Delta t,\quad\mathbf{p}^{(i)}_{t}\leftarrow\mathbf{p}^{(i)}_{t}$, (2)
$\text{(Low)}\;\mathbf{q}^{(i)}_{t}\leftarrow\mathbf{q}^{(i)}_{t},\quad\mathbf{p}^{(i)}_{t}\leftarrow\mathbf{p}^{(i)}_{t}+\beta(\mathbf{q}^{(i)}_{t})\Delta t.$
Here, we use $\alpha$ and $\beta$ that can be either linear or activation modules, representing learnable parameters, ensuring we remain in the class of canonical (i.e., symplectic) maps. By stacking multiple up/low blocks, we achieve a deep network $\Phi_{\theta}:(\mathbf{q^{(i)}_{t}},\mathbf{p^{(i)}_{t}},\Delta t)\mapsto(\mathbf{q^{(i)}_{t+1}},\mathbf{p^{(i)}_{t+1}})$. A crucial mathematical property is that the composition of symplectic transformations remains symplectic. Thus, when these modules are stacked together in sequence, regardless of network depth or configuration, the resulting transformation is guaranteed to be symplectic. This relies on the algebraic properties of the “Up” and “Low” updates, which act as symplectic shear transformations. For instance, the Jacobian matrix $\mathbf{J}_{low}$ of a “Low” update $(\mathbf{q},\mathbf{p})\mapsto(\mathbf{q},\mathbf{p}+\beta(\mathbf{q})\Delta t)$ is lower-triangular with identity diagonals:
$\mathbf{J}_{low}=\begin{pmatrix}\mathbf{I}&\mathbf{0}\\ \frac{\partial\beta}{\partial\mathbf{q}}\Delta t&\mathbf{I}\end{pmatrix}$. (3)
For a map to be symplectic, its Jacobian must satisfy $\mathbf{J}^{T}\bm{\Omega}\mathbf{J}=\bm{\Omega}$. This condition holds for triangular maps provided the cross-term block (here, $\frac{\partial\beta}{\partial\mathbf{q}}$) is a symmetric matrix. To ensure this, we implement $\alpha$ and $\beta$ either as gradients of a scalar potential or using layers with symmetric weights (Jin et al., 2020).
A pivotal feature of Hamiltonian dynamics is time reversibility, i.e., if we integrate the system forward from $(\mathbf{q^{(i)}_{t}},\mathbf{p^{(i)}_{t})}$ to $(\mathbf{q^{(i)}_{t+1}},\mathbf{p^{(i)}_{t+1}})$ over time $\Delta t$, then integrating backwards over $\Delta t$ should return the system exactly to $(\mathbf{q^{(i)}_{t}},\mathbf{p^{(i)}_{t}})$. This property lies at the heart of many physical invariants (energy, momentum, etc.) and is crucial for long-horizon stability in forecasting the behavior of such systems.
@@
During training the encoder is provided with a sequence of phase-space points (T time-steps) and encodes them onto a conservative latent space $\mathbf{z}_{c}=[\mathbf{q_{enc}},\mathbf{p_{enc}}]\in\mathbb{R}^{2d\times T}$ that is ultimately used by the ActiveDecoder in the cross attention mechanism to specifically fine-tune the query and value parameters.
Additionally, to ensure that our encoder can generalize quickly across multiple related but distinct physical systems, we adopt a MAML-style (Finn et al., 2017) framework. This meta-learning paradigm works particularly well in the SymplecticEncoder’s case, since SympNets represent canonical transformations that are known to have well-defined Hessian (Birtea et al., 2020), which is crucial for methods like MAML that use second-order gradients. For each system $i$ in a mini-batch, we split its trajectory into $\mathcal{I}_{adapt}$ and $\mathcal{I}_{meta}$ sets. During the fast adaptation loop, we optimize the parameters of the encoder $\theta_{SE}$ only on $\mathcal{I}_{adapt}$, by minimizing the mean-squared error between its forward predictions and the ground-truth labels. This process simulates “specializing” the encoder to system $i$’s local dynamics using a simple loss function that avoids gradient-terms that may destabilize the subsequent meta-update of $\theta_{SE}$. For the outer loop, we perform a forward and an inverse pass of the SymplecticEncoder and subsequently we minimize a combined loss represented as,
$\mathcal{L}_{\mathrm{meta}}=\frac{1}{\mathcal{T}_{meta}N_{batch}}$ $\sum_{\begin{subarray}{c}t\in\mathcal{I}_{\mathrm{meta}}\\ i\in N_{batch}\end{subarray}}\Bigl\|\Phi_{\theta_{SE}}\bigl(\mathbf{q}_{t}^{(i)},\mathbf{p}_{t}^{(i)};\bm{\theta_{SE}}^{*(i)}\bigr)-\Phi_{\theta_{SE}}^{-1}\bigl(\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)};\bm{\theta_{SE}}^{*(i)}\bigr)-[\mathbf{q}_{t+1}^{(i)}-\mathbf{q}_{t}^{(i)},\mathbf{p}_{t+1}^{(i)}-\mathbf{p}_{t}^{(i)}]\Bigr\|^{2},$ (4)
where $\mathcal{T}_{meta}$ is the total number of time-steps stored in $\mathcal{I}_{meta}$ set and $N_{batch}$ the total number of meta-learned systems in the batch. While the SymplecticEncoder is trained in isolation, we freeze its parameters during the ActiveDecoder’s training.